# Fourier restriction and well-approximable numbers

### v2

@@ 3.5. Fourier decay of $\mu$
Since $\operatorname*{supp}(\mu)\subseteq\mathbb{T}$, $\widehat{\mu}(\xi)=\mathcal{F}(\mu)(\xi)$ for integer $\xi$. So the preceding proposition shows that $\mu$ satisfies (1.2) and (1.5) when $\xi\in\mathbb{Z}$. In fact, by a standard argument (see, e.g., [20, pp.252-253] or [34, p.69]) the same estimates extend to all $\xi\in\mathbb{R}$. Thus we have proved $\mu$ satisfies (1.2) if $\beta\geq 0$ and (1.5) if $\beta<0$.
@@ 3.6. Regularity of $\mu$
$\begin{split}g_{k}F_{I}(x)=&\frac{M_{k}^{1+\alpha}}{|P_{M}|+M_{k}^{\beta}+1}\sum_{(p,r)\in P_{M_{k}}(I)}F(M_{k}^{1+\alpha}(px-r))F_{I}(x)\\ &+\frac{M_{k}^{1+\alpha}(1+M_{k}^{-\beta})}{|P_{M}|+M_{k}^{\beta}+1}\sum_{a\in B_{k}(I)}F(M_{k}^{\alpha-\beta+1}(B_{k}x-a))F_{I}(x).\end{split}$
@@
$g_{k}(x)\leq\log^{2}(M_{k})M_{k}^{\alpha}\leq M_{k}^{1+\alpha}.$
@@
$\begin{split}g_{1}\cdots g_{k}(x)&\leq(M_{1}M_{2}\cdots M_{k-1})^{1+\alpha}\log^{2}(M_{k})M_{k}^{\alpha}\\ &\leq\log(M_{2})M_{2}^{1+\alpha}(M_{3}\cdots M_{k-1})^{1+\alpha}\log^{2}(M_{k})M_{k}^{\alpha}\\ &\leq\log(M_{3})M_{3}^{1+\alpha}(M_{4}\cdots M_{k-1})^{1+\alpha}\log^{2}(M_{k})M_{k}^{\alpha}\\ &\qquad\vdots\\ &\leq\log^{3}(M_{k})M_{k}^{\alpha}.\end{split}$
@@
Also, by (3.23) we have that
@@
$\begin{split}\sum_{J\in\mathcal{J}_{1}}\mu(J)&\lesssim\log^{3}(M_{k-1})M_{k-1}^{\alpha}|P_{M_{k}}|\log(M_{k})M_{k}^{-1}|I|\\ &\lesssim\log^{3}(M_{k-1})M_{k-1}^{\alpha}|I|.\end{split}$ (3.30)
@@ 4. Failure of the extension estimate
In this section, we define a sequence of function $f_{k}$ which witness the failure of the estimate (R). More specifically, we shall prove (1.3) if $\beta\geq 0$ and (1.6) if $\beta<0$. This will complete the proofs of Theorem 1.7 and Theorem 1.8.
@@ 4.1. Definitions and Lemmas
If $10<\beta<$ 1,
$|\widehat{h}_{2}(s)|\lesssim\frac{\log(M_{k})}{M_{k}}\left(1+\frac{|s|}{M_{k}^{2+\alpha}}\right)^{-10}.$ (4.1)
If $-1<\beta<0$,
$|\widehat{h}_{2}(s)|\lesssim\frac{1}{M_{k}}\left(1+\frac{|s|}{M_{k}^{2+\alpha}}\right)^{-10}.$ (4.2)
@@
First, let us consider when $10<\beta<$ 1. Note that there are at most $\frac{\log(|s-t|)}{\log M_{k}}$ values of $p\in P_{M_{k}}$ such that $p|(s-t)$. For fixed $s$, if $|s-t|\geq|s|/2$,
$\begin{split}\left|\sum_{p\in P_{M_{k}}}\widehat{F}\left(\frac{s-t}{pM_{k}^{1+\alpha}}\right)\mathbf{1}_{p\mathbb{Z}}(s-t)\right|&\lesssim\frac{\log(|s-t|)}{\log M_{k}}\left(1+\frac{|s-t|}{M_{k}^{2+\alpha}}\right)^{-11}\\ &\lesssim\left(1+\frac{|s|}{M_{k}^{2+\alpha}}\right)^{-10}.\end{split}$
If $|s-t|\leq|s|/2$, then $|s|\sim|t|$. Thus,
$\left|\widehat{F}\left(\frac{t}{B_{k}M_{k}^{1+\alpha-\beta}}\right)\right|\lesssim\left(1+\frac{|s|}{M_{k}^{2+\alpha}}\right)^{-10}\left(1+\frac{|t|}{M_{k}^{2+\alpha}}\right)^{-10}$
and
$\begin{split}\left|\sum_{p\in P_{M_{k}}}\widehat{F}\left(\frac{s-t}{pM_{k}^{1+\alpha}}\right)\mathbf{1}_{p\mathbb{Z}}(s-t)\right|&\lesssim\frac{\log(|s-t|)}{\log M_{k}}\left(1+\frac{|s-t|}{M_{k}^{2+\alpha}}\right)^{-1}\\ &\lesssim 1.\end{split}$
Therefore, we have
$|\widehat{h}_{2}(s)|\lesssim\frac{M_{k}^{\beta-\alpha-1}}{|P_{M_{k}}|}\left(1+\frac{|s|}{M_{k}^{2+\alpha}}\right)^{-10}\sum_{t\in\mathbb{Z}}\left(1+\frac{|t|}{M_{k}^{2+\alpha}}\right)^{-10}\mathbf{1}_{B_{k}\mathbb{Z}}(t).$
If $\mathbf{1}_{B_{k}\mathbb{Z}}(t)\neq 0$, then $t=B_{k}m$ for some $m\in\mathbb{Z}$. Thus,
$\begin{split}|\widehat{h}_{2}(s)|&\lesssim\frac{M_{k}^{\beta-\alpha-1}}{|P_{M_{k}}|}\left(1+\frac{|s|}{M_{k}^{2+\alpha}}\right)^{-10}\sum_{m\in\mathbb{Z}}\left(1+\frac{|B_{k}m|}{M_{k}^{2+\alpha}}\right)^{-10}\\ &\lesssim\frac{\log(M_{k})}{M_{k}}\left(1+\frac{|s|}{M_{k}^{2+\alpha}}\right)^{-10}.\end{split}$
Thus, we obtain (4.1)
Let us consider when $-1<\beta<0$. For fixed $s$, if $|s-t|\geq|s|/2$,
$\left|\widehat{F}\left(\frac{s-t}{B_{k}M_{k}^{1+\alpha-\beta}}\right)\right|\lesssim\left(1+\frac{|s|}{M_{k}^{2+\alpha}}\right)^{-10}.$
@@
Therefore, we have
@@
$\begin{split}|\widehat{h}_{2}(s)|&\lesssim\frac{M_{k}^{\beta-\alpha-1}}{|P_{M_{k}}|}\left(1+\frac{|s|}{M_{k}^{2+\alpha}}\right)^{-10}\sum_{p\in P_{M_{k}}}\sum_{m\in\mathbb{Z}}\left(1+\frac{|pB_{k}m+B_{k}a_{s,p}|}{M_{k}^{2+\alpha}}\right)^{-10}\\ &\lesssim\frac{M_{k}^{\beta-\alpha-1}}{|P_{M_{k}}|}\left(1+\frac{|s|}{M_{k}^{2+\alpha}}\right)^{-10}\sum_{p\in P_{M_{k}}}\sum_{m\in\mathbb{Z}}\left(1+\frac{|m+a_{s,p}/p|}{M_{k}^{\alpha-\beta}}\right)^{-10}\\ &\lesssim\frac{1}{M_{k}}\left(1+\frac{|s|}{M_{k}^{2+\alpha}}\right)^{-10}.\end{split}$
In the last inequality, we used that $\beta<0<\alpha$. Therefore, we obtain (4.2).
@@
Let $l>k$. For $|m|\leq M_{k}^{1+\alpha-\beta}/10$, we have
@@
We claim that the following estimates hold for sufficiently large $k$ and for $|m|\leq M_{k}^{1+\alpha-\beta}/10$:
@@
$\sum_{t\in\mathbb{Z}}|\widehat{h}_{2}(B_{k}m-t)(g_{l}\cdots g_{k+1}\mu_{k-1})^{\wedge}(t)|\lesssim\log^{2}(M_{k})M_{k}^{-1}\qquad\text{if}\ 0<\beta<1;$ (4.6)
$\sum_{t\in\mathbb{Z}}|\widehat{h}_{2}(B_{k}m-t)(g_{l}\cdots g_{k+1}\mu_{k-1})^{\wedge}(t)|\lesssim\log^{1/10}(M_{k})M_{k}^{-1}\qquad\text{if}\ -1<\beta<0.$ (4.7)
By combining (4.4), (4.5), (4.6) and (4.7) we get (4.3) when $\beta\neq 0$.
@@
Since $F$ is a Schwartz function such that $F(x)\gtrsim 1$ if $|x|\leq 1/4$, we have
@@
Using $|\widehat{h}_{1}(s)|\lesssim\log(M_{k})M_{k}^{-1}(1+M_{k}^{\beta})$, we obtain
@@
Proof of (4.6) and (4.7)
Let us consider when $10<\beta<$ 1. By the same method we used in the proof of (4.5), we use Lemma 4.1 and obtain
$\sum_{|t|\geq M_{k}^{10(2+\alpha)}}|\widehat{h}_{2}(B_{k}m-t)(g_{l}\cdots g_{k+1}\mu_{k-1})^{\wedge}(t)|\lesssim\log(M_{k})M_{k}^{-1-80(2+\alpha)}.$
By (3.2), we have $\log(M_{k})^{1/10}\geq 2M_{k-1}^{2+\alpha}$. By Lemma 4.1, $|\widehat{h}_{2}(s)|\lesssim\log(M_{k})M_{k}^{-1}$. Thus, we use (3.10) with $N=99$ and (3.19) to obtain
$\begin{split}\sum_{\log^{1/10}(M_{k})\leq|t|\leq M_{k}^{10(2+\alpha)}}&|\widehat{h}_{2}(B_{k}m-t)(g_{l}\cdots g_{k+1}\mu_{k-1})^{\wedge}(t)|\\ &\lesssim\log(M_{k})M_{k}^{-1}\left(M_{k}^{10(2+\alpha)}M_{k+1}^{-90}+\log(M_{k})\sum_{|t|\geq\log^{1/10}(M_{k})}\log(|t|)|t|^{-99}\right)\\ &\lesssim\left(\log^{5}(M_{k})M_{k}\right)^{-1}.\end{split}$
@@
$\sum_{|t|\leq\log(M_{k})^{1/10}}|\widehat{h}_{2}(B_{k}m-t)(g_{l}\cdots g_{k+1}\mu_{k-1})^{\wedge}(t)|\lesssim\log^{2}(M_{k})M_{k}^{-1}.$ (4.9)
@@
When $-1<\beta<0$, we the proof is same except that $\log(M_{k})M_{k}^{-1}$ is replaced by $M_{k}^{-1}$ because of Lemma 4.1. Also, instead of (4.9), we have
$\sum_{|t|\leq\log(M_{k})^{1/10}}|\widehat{h}_{2}(B_{k}m-t)(g_{l}\cdots g_{k+1}\mu_{k-1})^{\wedge}(t)|\lesssim\log^{1/10}(M_{k})M_{k}^{-1}.$
Therefore, we obtain (4.7).
@@
In the last inequality, we used (4.10). The bound on $\xi$ guarantees that this is much smaller than $\log(M_{k})M_{k}^{\beta-1}$. By Lemma 4.2, for large $k$, we have that $\mathcal{F}(f_{k}\mu)(B_{k}m+\xi)\gtrsim\log(M_{k})M_{k}^{\beta-1}$ whenever $|m|\leq M_{k}^{1+\alpha-\beta}/10$ Thus, we obtain

### v3

@@ 3.1. Preliminaries
Because $F$ is a Schwartz function, so is $\mathcal{F}(F^{2})$. Since $F^{2}$ has a positive integral, $\mathcal{F}({F^{2}})(0)>0$. We can therefore choose a constant $C_{F}$ such that $\mathcal{F}({F^{2}})(\xi)\gtrsim 1$ provided that $|\xi|\leq C_{F}$.
@@ 3.5. Fourier decay of $\mu$
Since $\operatorname*{supp}(\mu)\subseteq\mathbb{T}$, $\widehat{\mu}(\xi)=\mathcal{F}(\mu)(\xi)$ for integer $\xi$. So the preceding proposition shows that $\mu$ satisfies (1.2) and (1.5) when $\xi\in\mathbb{Z}$. In fact, by a standard argument (see, e.g., [20, p.252-253] or [34, p.69]) the same estimates extend to all $\xi\in\mathbb{R}$. Thus we have proved $\mu$ satisfies (1.2) if $\beta\geq 0$ and (1.5) if $\beta<0$.
@@ 3.6. Regularity of $\mu$
$\begin{split}g_{k}F_{I}(x)=&\frac{M_{k}^{1+\alpha}}{|P_{M_{k}}|+M_{k}^{\beta}+1}\sum_{(p,r)\in P_{M_{k}}(I)}F(M_{k}^{1+\alpha}(px-r))F_{I}(x)\\ &+\frac{M_{k}^{1+\alpha}(1+M_{k}^{-\beta})}{|P_{M_{k}}|+M_{k}^{\beta}+1}\sum_{a\in B_{k}(I)}F(M_{k}^{\alpha-\beta+1}(B_{k}x-a))F_{I}(x).\end{split}$
@@
$g_{k}(x)\leq\log^{2}(M_{k})M_{k}^{\alpha}\leq M_{k}^{1+\alpha},$
and for $11\leq j\leq k-$ 1, we have the weaker estimate
$g_{j}(x)\leq C_{0}D\log(M_{j})M_{j}^{\alpha}(2+M_{j}^{-\beta})\leq\log^{2}(M_{j})M_{j}^{\alpha}(2+M_{j}^{-\beta}).$
@@
$\begin{split}g_{1}\cdots g_{k}(x)&\leq\left(\prod_{j=1}^{k-1}\log^{2}(M_{j})M_{j}^{\alpha}(2+M_{j}^{-\beta})\right)\log^{2}(M_{k})M_{k}^{\alpha}\\ &\leq\log(M_{2})\left(\prod_{j=2}^{k-1}\log^{2}(M_{j})M_{j}^{\alpha}(2+M_{j}^{-\beta})\right)\log^{2}(M_{k})M_{k}^{\alpha}\\ &\leq\log(M_{3})\left(\prod_{j=3}^{k}\log^{2}(M_{j})M_{j}^{\alpha}(2+M_{j}^{-\beta})\right)\log^{2}(M_{k})M_{k}^{\alpha}\\ &\qquad\vdots\\ &\leq\log^{3}(M_{k})M_{k}^{\alpha}.\end{split}$
@@
Also, by truncating the above argument one step earlier and using (3.23), we have that
@@
$\begin{split}\sum_{J\in\mathcal{J}_{1}}\mu(J)&\lesssim\log^{3}(M_{k-1})M_{k-1}^{\alpha}(2+M_{k-1}^{-\beta})|P_{M_{k}}|\log(M_{k})M_{k}^{-1}|I|\\ &\lesssim\log^{3}(M_{k-1})M_{k-1}^{\alpha}(2+M_{k-1}^{-\beta})|I|.\end{split}$ (3.30)
@@
$\sum_{J\in\mathcal{J}_{1}}\mu(J)\lesssim\log^{3}(|I|^{-1})(|I|^{\frac{2}{2+\alpha}}+|I|^{\frac{2+\beta}{2+\alpha}}).$
@@ 4. Failure of the extension estimate
In this section, we define a sequence of functions $f_{k}$ which witness the failure of the estimate (R). More specifically, we shall prove (1.3) if $\beta\geq 0$ and (1.6) if $\beta<0$. This will complete the proofs of Theorem 1.7 and Theorem 1.8.
@@ 4.1. Definitions and Lemmas
If $\beta\neq 0$,
$|\widehat{h}_{2}(s)|\lesssim\frac{1+M_{k}^{\beta-\alpha}}{M_{k}}\left(1+\frac{|s|}{M_{k}^{2+\alpha}}\right)^{-10}.$ (4.1)
@@
For fixed $s$, if $|s-t|\geq|s|/2$,
$\left|\widehat{F}\left(\frac{s-t}{pM_{k}^{1+\alpha}}\right)\right|\lesssim\left(1+\frac{|s|}{M_{k}^{2+\alpha}}\right)^{-10},$
and because $F$ is a Schwartz function, we also have that
$\left|\widehat{F}\left(\frac{t}{B_{k}M_{k}^{1+\alpha-\beta}}\right)\right|\lesssim\left(1+\frac{|t|}{M_{k}^{2+\alpha}}\right)^{-10}.$
@@
and we have the trivial bound
$\left|\widehat{F}\left(\frac{s-t}{pM_{k}^{1+\alpha}}\right)\right|\lesssim 1.$
Therefore, we have for all $t\in\mathbb{Z}$ that
$\left|\widehat{F}\left(\frac{t}{B_{k}M_{k}^{1+\alpha-\beta}}\right)\widehat{F}\left(\frac{s-t}{pM_{k}^{1+\alpha}}\right)\right|\lesssim\left(1+\frac{|s|}{M_{k}^{2+\alpha}}\right)^{-10}\left(1+\frac{|t|}{M_{k}^{2+\alpha}}\right)^{-10}.$
Thus,
@@
$\begin{split}|\widehat{h}_{2}(s)|&\lesssim\frac{M_{k}^{\beta-\alpha-1}}{|P_{M_{k}}|}\left(1+\frac{|s|}{M_{k}^{2+\alpha}}\right)^{-10}\sum_{p\in P_{M_{k}}}\sum_{m\in\mathbb{Z}}\left(1+\frac{|pB_{k}m+B_{k}a_{s,p}|}{M_{k}^{2+\alpha}}\right)^{-10}\\ &\lesssim\frac{M_{k}^{\beta-\alpha-1}}{|P_{M_{k}}|}\left(1+\frac{|s|}{M_{k}^{2+\alpha}}\right)^{-10}\sum_{p\in P_{M_{k}}}\sum_{m\in\mathbb{Z}}\left(1+\frac{|m+a_{s,p}/p|}{M_{k}^{\alpha-\beta}}\right)^{-10}.\end{split}$
It is easy to show that
$\sum_{m\in\mathbb{Z}}\left(1+\frac{|m+a_{s,p}/p|}{M_{k}^{\alpha-\beta}}\right)^{-10}\lesssim\begin{cases}M_{k}^{\alpha-\beta}&\text{if }\beta\leq\alpha,\\ 1&\text{if }\alpha<\beta.\end{cases}$
Therefore, we obtain
$|\widehat{h}_{2}(s)|\lesssim\frac{1+M_{k}^{\beta-\alpha}}{M_{k}}\left(1+\frac{|s|}{M_{k}^{2+\alpha}}\right)^{-10}.$
@@
Let $l>k$. For $|m|\leq M_{k}^{1+\alpha-\beta}/C_{F}$, we have
@@
We claim that the following estimates hold for sufficiently large $k$ and for $|m|\leq M_{k}^{1+\alpha-\beta}/C_{F}$:
@@
$\sum_{t\in\mathbb{Z}}|\widehat{h}_{2}(B_{k}m-t)(g_{l}\cdots g_{k+1}\mu_{k-1})^{\wedge}(t)|\lesssim\log^{1/10}(M_{k})M_{k}^{-1}(1+M_{k}^{\beta-\alpha})\qquad\text{if}\ \beta\neq 0.$ (4.5)
Recall that $\alpha>0$. By combining (4.3), (4.4) and (4.5), we get (4.2) when $\beta\neq 0$.
@@
By our choice of the constant $C_{F}$, the assumption $|m|\leq M_{k}^{1+\alpha-\beta}/C_{F}$ implies
@@
Using $|\widehat{h}_{1}(s)|\lesssim\log(M_{k})M_{k}^{-1}(1+M_{k}^{\beta})$ and (3.19), we obtain
@@
Proof of (4.5)
By the same method we used in the proof of (4.4), we use Lemma 4.1 and obtain
$\sum_{|t|\geq M_{k}^{10(2+\alpha)}}|\widehat{h}_{2}(B_{k}m-t)(g_{l}\cdots g_{k+1}\mu_{k-1})^{\wedge}(t)|\lesssim(1+M_{k}^{\beta-\alpha})M_{k}^{-1-80(2+\alpha)}.$
By (3.2), we have $\log(M_{k})^{1/10}\geq 2M_{k-1}^{2+\alpha}$. By Lemma 4.1, $|\widehat{h}_{2}(s)|\lesssim M_{k}^{-1}(1+M_{k}^{\beta-\alpha})$. Thus, we use (3.10) with $N=99$ and (3.19) to obtain
$\begin{split}\sum_{\log^{1/10}(M_{k})\leq|t|\leq M_{k}^{10(2+\alpha)}}&|\widehat{h}_{2}(B_{k}m-t)(g_{l}\cdots g_{k+1}\mu_{k-1})^{\wedge}(t)|\\ &\lesssim M_{k}^{-1}(1+M_{k}^{\beta-\alpha})\left(M_{k}^{10(2+\alpha)}M_{k+1}^{-90}+\log(M_{k})\sum_{|t|\geq\log^{1/10}(M_{k})}\log(|t|)|t|^{-99}\right)\\ &\lesssim\log^{-5}(M_{k})M_{k}^{-1}(1+M_{k}^{\beta-\alpha}).\end{split}$
@@
$\sum_{|t|\leq\log(M_{k})^{1/10}}|\widehat{h}_{2}(B_{k}m-t)(g_{l}\cdots g_{k+1}\mu_{k-1})^{\wedge}(t)|\lesssim\log^{1/10}(M_{k})M_{k}^{-1}(1+M_{k}^{\beta-\alpha}).$ (4.7)
@@
In the last inequality, we used (4.8). The bound on $\xi$ guarantees that this is much smaller than $\log(M_{k})M_{k}^{\beta-1}$. By Lemma 4.2, for large $k$, we have that $\mathcal{F}(f_{k}\mu)(B_{k}m+\xi)\gtrsim\log(M_{k})M_{k}^{\beta-1}$ whenever $|m|\leq M_{k}^{1+\alpha-\beta}/C_{F}$ Thus, we obtain